# OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion

### v1

@@ 4 Results
Our primary motivation for the proposed approach is to build an E2E ST system that is context-aware, efficient for SimulST and enabled to exploit additional image context. To validate this, we first evaluate the system in simultaneous scenarios by comparing latency and quality. Next, we report quality in offline scenarios and conduct an error analysis to contrast the E2E system with traditional cascaded systems fine-tuned on the same data. We then assess performance in TIT scenarios, benchmarking against other off-the-shelf multimodal translation models. Finally, we perform an ablation study to analyze the contributions of individual layers in the MMFM and present the key findings.
We also validate the proposed approach for a different model combination (Qwen-Audio (Chu et al., 2023) and Tower-Instruct (Alves et al., 2024)), analyze training with different layers and post-editing in our preliminary experiments and report the results in the Appendixes Table˜4 and table˜5.
@@ 4.1 Simultaneous Speech Translation
A SimulST system must generate translations incrementally as the source audio arrives, rather than waiting for the entire utterance to finish. The decision of whether to commit a token or read more audio is referred to as the policy. Since our E2E system does not natively support simultaneous translation, we adopt the widely used Local Agreement policy (Liu et al., 2020), which commits the longest common prefix observed across consecutive predictions. Concretely, at each step we process the audio in fixed chunk sizes (a hyperparameter) and irreversibly emit the overlapping prefix between the current and previous hypotheses. After committing, subsequent tokens are force-decoded conditioned on the partial translation committed so far.
To fairly evaluate the proposed approach, we compare our E2E model OmniFusion against a strong cascaded counterpart built from the same pretrained components. We therefore fine-tune this cascaded baseline on the same data using LoRA adapters (Omni + SeedX LoRA FT). The effect of fine-tuning is discussed in more detail in Section˜4.2. We experiment with chunk sizes of $[1,1.5,2,2.5,3]$ seconds, run inference both with and without images, and report computation-aware and computation-unaware AL together with XCOMET-XL scores for the $en\rightarrow de$ and $en\rightarrow it$ language pairs in Figure˜2.
The results reveal several key insights. First, as expected, the offline ST systems (with images) achieve the highest quality in both language pairs due to access to full context. Next, OmniFusion is consistently around 1 second faster than its cascaded counterpart across all chunk sizes while preserving comparable quality in the audio-only setting. When images are provided, OmniFusion delivers both the lowest latency and the highest quality, demonstrating that visual context is substantially more beneficial in the E2E model than in the cascaded baseline allowing it to commit high quality stable outputs earlier. Finally, we also see that in computation-aware settings, adding image causes delay due to processing and causes overhead than only-audio creating a trade-off between quality and latency.
We attribute the advantage of OmniFusion to two factors: (1) our fine-tuning strategy, which includes an OCR objective that tightly aligns image and audio representations, and (2) the inherent efficiency of the E2E architecture, which requires fewer decoding steps overall. In contrast, the cascaded pipeline with Local Agreement must wait for stable ASR hypotheses before translation can even begin, making it slower than the E2E approach.
In conclusion, OmniFusion achieves better latency, leverages image context effectively, and has higher quality in SimulST scenarios, making it particularly well-suited for real-world applications such as live presentations and meetings where visual information is readily available.
@@ 4.2 Offline Speech Translation
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 85.88 1212.3 773.0 74.3 2059.7
@@
End-to-End (Direct ST)
OmniFusion
Mid Fusion 83.57 1114.0 1188.3 72.3 2374.7
@@
Gated Fusion + Self-Casade 86.57 1231.0 719.0 55.3 2005.3
@@
Although SimulST is the primary focus of this work, we also evaluate the systems in the offline scenario, where latency is not a concern and the goal is to maximize translation quality. In this setting, we can leverage the self-cascade inference mode described in Section˜2.2 to further boost performance.
To measure the contributions of fine-tuning and visual context, we first evaluate the cascaded baseline in four configurations: base vs. fine-tuned (LoRA FT) models, each tested with and without images. For the E2E OmniFusion model, we additionally train a variant that uses only the middle-layer fusion without gating to assess the need for the gating module. We then compare all E2E configurations (mid-only fusion and full gated fusion) both with and without images, and with/without self-cascade inference. Results are averaged across three target languages $en\rightarrow\{de,it,zh\}$ and reported in Table˜2 using XCOMET-XL scores. We further include the detailed XCOMET-XL error categories, hypothesizing that E2E models may sacrifice minor fluency for substantially fewer major and critical errors.
The results yield several insights. In the cascaded pipeline, LoRA fine-tuning yields consistent gains (e.g., 85.94 $\rightarrow$ 86.59 with images), whereas adding images provides only marginal benefit. In contrast, the E2E OmniFusion model benefits substantially from both gated fusion (outperforming mid-only fusion) and image context showing the need for gating module and additional contextual information. Although the direct E2E translation score remains slightly below the best cascaded system, enabling self-cascade inference closes this gap almost entirely (cascaded LoRA FT: 86.59 vs. E2E gated fusion + self-cascade: 86.57). More importantly, error analysis reveals that our best E2E configuration produces the fewest critical and major errors (e.g., 719 vs. 751.3 for the strongest cascaded baseline). These findings demonstrate that OmniFusion, when equipped with image context and self-cascade inference, matches fine-tuned cascaded system in offline ST quality while offering significantly fewer severe translation errors
4.3 Text-Image Translation
@@
ZeroMMT 3.3B
(Futeral et al., 2025)
- 80.64 86.69 83.54 83.40 83.87 77.85
TowerVision 9B
(Viveiros et al., 2025)
- 70.85 86.62 84.97 84.29 85.81 85.24
@@
Table 3: COMET scores ($\uparrow$) on the CoMMuTE dataset. Statistically significant top-performing results per language pair are highlighted in bold. ‘Config‘ distinguishes the two OmniFusion variants (mid-layer fusion vs. gated fusion).
Apart from ST, translating captions also relies on image context, where cascaded systems fail due to their inability for disambiguation. To demonstrate this, we evaluate several models on the CoMMuTE test set across multiple language pairs and report COMET scores in Table˜3.
@@
First, the audio-only Omni 2.5B exhibits the weakest multilingual performance, which reinforces our motivation for integrating it with a strong translation backbone. SeedX 7B obtains reasonable scores yet consistently trails all vision-capable models, confirming its limitation in the absence of image context. Among vision-aware systems, both Mid Fusion and Gated Fusion variants of our OmniFusion 14B achieve state-of-the-art or near-SOTA COMET scores, matching or surpassing vision-only models such as ZeroMMT 3.3B (Futeral et al., 2025) and TowerVision 9B (Viveiros et al., 2025) on most language directions.
Interestingly, and in contrast to ST, the Mid Fusion variant slightly outperforms Gated Fusion on CoMMuTE. This suggests that, when the image serves primarily as supplementary context (as in caption translation) rather than the main information source (as in audio for spoken ST), using visual features at the middle layer alone may be sufficient, and the additional flexibility of gating offers no further advantage.
Importantly, unlike TowerVision which requires massive multimodal pretraining from scratch, OmniFusion can be built on top of existing strong pretrained components. This makes our approach far more data- and training-efficient without compromising quality, by effectively exploiting the rich image–text alignment knowledge already present in the MMFM pretrained weights. These results underline the strength of our E2E design on visually grounded translation tasks where image context is essential.
4.4 Contribution of Layers
@@
We presented a gating module in Section˜2.1 for fusing information across layers of the MMFM. To better understand its behavior, we conduct an ablation study by examining the gate values assigned to each MMFM layer during fusion. Specifically, we compute the average gate value across the three layers (first, middle, last), noting that the gate weights for each timestep sum to 1.
We run inference on 100 examples from the MCIF dataset—both speech-only and speech+image conditions—and on the CoMMuTE dataset for the image-only modality. The resulting average gate values are shown in Figure˜3.
We observe that the first and middle layers receive consistently higher contributions, while the last layer contributes only marginally. This suggests that the majority of transferable multimodal processing occurs in the earlier half of the MMFM. Although the weak contribution of the last layer is not entirely clear, we hypothesize that these representations may be more language-specific, whereas the middle layers capture more language-agnostic features (Liu and Niehues, 2025).
@@ 5 Related Work
Multimodal MT: Translating captions has been widely explored in the MT field over the last decade (Specia et al., 2016; Elliott et al., 2017; Barrault et al., 2018). Several works focus on enabling NMT models to process images, either by training models from scratch (Vijayan et al., 2024) or by augmenting large pretrained models such as NLLB (Costa-Jussà et al., 2022; Futeral et al., 2025). Most recently, and in parallel to our work, Viveiros et al. (2025) extend Tower with image and video modalities through a computationally expensive multi-stage training pipeline.
While these efforts focus primarily on vision, an other line of direction is adding audio to translation LLMs (Ambilduke et al., 2025). A smaller line of work tackles the combination of vision and audio, such as Sinhamahapatra and Niehues (2025), which exploits slides and speech for ASR. Building on this trend, there is also increasing interest in translating audio together with video (Anwar et al., 2023; Choi et al., 2024). In contrast to these works, our approach focuses on jointly using images and audio specifically for ST, while remaining efficient in both training and inference.
Simultaneous ST: Many real-world scenarios require real-time ST without waiting for the full source utterance. However, most research focuses on training models in offline settings and only later adapting them to SimulST. Such adaptation typically relies on predefined or learned translation policies, such as wait-k (Ma et al., 2019; Elbayad et al., 2020), which assumes a fixed token delay before committing to output. Other approaches include Local Agreement (Liu et al., 2020), which selects the most stable prefix across consecutive read steps or attention-based heuristics (Papi et al., 2023). More recently, several works explore leveraging LLMs for SimulST policy learning (Koshkin et al., 2024b, a; Guo et al., 2025; Fu et al., 2025). In our work, we use Local Agreement to validate the effectiveness of our approach and to compare against cascaded baselines. However, our method is compatible in theory with other SimulST policies, and integrating more advanced or LLM-driven strategies is a promising direction for future work.
Modality Fusion: With the rapid development of multimodal foundation models, a key challenge is how to effectively integrate speech and image modalities into text-centric LLMs. Some approaches (Abdin et al., 2024; Xu et al., 2025; Team et al., 2025) process each modality with separate encoders and merge them later within the LLM backbone. In contrast, models such as (Ye et al., 2025) enable early interaction between vision and audio modalities, facilitating better cross-modal alignment. Beyond the timing of modality interaction, ongoing research also investigates the design of modality projectors (Verdini et al., 2024; Gaido et al., 2024b) and the use of continuous versus discrete tokens for non-text modalities (Zhan et al., 2024; Li et al., 2025), balancing the trade-offs between latency and translation quality.
@@ 6 Conclusion
We proposed a novel approach to fuse a MMFM with a translation LLM, and validated its effectiveness across three key scenarios. First, we showed that the fused model achieves faster inference while maintaining strong translation quality in SimulST settings, even when only partial source audio is available. Second, the model produces fewer major and critical translation errors, demonstrating more reliable linguistic accuracy. Third, we found that the model effectively leverages visual context for both speech and text translation, leading to consistent gains in multimodal settings. Finally, our layer-level analysis reveals that the first and middle MMFM layers contribute most to capturing semantic multimodal information, providing insight into where meaningful cross-modal representations emerge.
For future work, our fusion framework can naturally be extended to additional modalities. Since Omni already supports video processing, incorporating video embeddings would enable richer context for translation. Moreover, the fusion strategy is not limited to models for multimodal inputs but future extensions could target multimodal generation, allowing the system to produce not only text but also speech, images, or video as output.
@@
A primary limitation of our work is that OmniFusion is trained with fixed task-specific prompts across the three supported translation settings, and therefore does not exhibit general instruction-following capabilities. It remains unclear whether our fusion strategy can be extended to support more flexible, instruction-driven multimodal behavior. In addition, most of our experiments use English as the source language. Although we report zero-shot results in Table 4, we do not evaluate the OmniFusion model with vision inputs in multilingual zero-shot conditions.

### v2

@@ 4 Results
Our goal is to develop a high-quality multimodal translation system with low-latency end-to-end SimulST that effectively leverages image context. We evaluate latency–quality trade-offs in simultaneous settings, compare offline translation quality and errors against cascaded models trained on the same data, assess text-and-image translation performance against multimodal baselines, and present ablations on MMFM layers and fusion mechanisms.
We also report ablations on training strategies to disentangle architectural and training effects (Appendix Table˜9). Additionally, we include preliminary experiments with alternative model pairings and layer selections (Qwen-Audio (Chu et al., 2023) and Tower-Instruct (Alves et al., 2024)) as well as post-editing results (Appendix Tables˜5 and 8). We observe similar findings and do not explore additional model combinations due to computational constraints.
@@ 4.1 Simultaneous Speech Translation
A SimulST system must translate incrementally as audio arrives, requiring a policy to decide when to read or commit output. Since our E2E model does not natively support simultaneous decoding, we adopt the Local Agreement policy (Liu et al., 2020), which commits the longest common prefix between consecutive hypotheses. Audio is processed in fixed-size chunks, and committed tokens are force-decoded in subsequent steps.
For a fair comparison, we evaluate OmniFusion against a strong cascaded baseline built from the same pretrained components and fine-tuned on the same data using LoRA adapters (Omni + SeedX). We experiment with chunk sizes of $[1,1.5,2,2.5,3]$ seconds, with and without images. As computational time in SimulST is critical, we report both computation-aware and -unaware Average Lagging along with XCOMET-XL scores for $en\rightarrow de$ and $en\rightarrow it$ in Figure˜3.
Offline ST with images yields the highest quality, as expected. In simultaneous settings, OmniFusion is consistently about 1 second faster than the cascaded system across all chunk sizes, with comparable or slightly lower quality in the audio-only condition. When images are provided, OmniFusion achieves both the lowest latency and highest quality, indicating that visual context is more effectively exploited in the E2E model, enabling earlier and more stable commitments. Image processing introduces additional delay, highlighting a latency–quality trade-off. We hypothesize that OmniFusion’s advantage stems from (i) our fine-tuning strategy, including an OCR objective that aligns audio and visual representations (Table˜9 shows performance degradations without multitask objectives) and (ii) the efficiency of the end-to-end architecture, which prevents the delayed onset of translation inherent in cascaded pipelines. Combined with Local Agreement decoding, OmniFusion achieves improved latency–quality trade-offs compared to cascaded systems.
@@ 4.2 Offline Speech Translation
Omni FT (Direct ST) LoRA FT 82.30 1079.0 852.0 90.3 2021.3
Cascade Systems (ASR → MT) Omni + SeedX No FT 85.88 1212.3 773.0 74.3 2059.7
@@
End-to-End (Direct ST) OmniFusion Mid Fusion 83.57 1114.0 1188.3 72.3 2374.7
@@
Gated Fusion + Self-Cascade 86.57 1231.0 719.0 55.3 2005.3
@@
ZeroMMT 3.3B - 80.64 86.69 83.54 83.40 83.87 77.85
TowerVision 9B - 70.85 86.62 84.97 84.29 85.81 85.24
@@
Table 2: COMET scores ($\uparrow$) on the CoMMuTE dataset. Statistically significant top-performing results per language pair are highlighted in bold. Qualitative example can be found in Table˜16
Although SimulST is our primary focus, we also evaluate all systems in an offline setting where latency is unconstrained and translation quality is maximized. In this scenario, we additionally leverage the self-cascade inference mode (Section˜2.3) to further improve performance. To isolate the effects of fine-tuning and visual context, we evaluate the cascaded baseline in five configurations: base MMFM fine-tuned for ST [^7] Even if it is competitive, it is insufficient due to limited language coverage compared to SeedX, and base vs. LoRA cascade fine-tuned with and without images. For OmniFusion, we train an additional mid-only fusion variant (without gating) (We find using the middle layer performs best in our preliminary experiments from Table 5) to assess the contribution of the gating module. We compare all E2E configurations with and without images and self-cascade inference. Results are averaged over $en\rightarrow{de,it,zh}$ and reported using XCOMET-XL in Table˜1, including fine-grained error categories. Additionally, inference times are reported in Table 7, although latency is not a primary constraint.
First, directly fine-tuning the MMFM for speech translation performs substantially worse than OmniFusion (82.30 vs 83.98 XCOMET), confirming combining is more effective than adapting the MMFM alone. In the cascaded pipeline, LoRA fine-tuning provides consistent gains (e.g., 85.94 $\rightarrow$ 86.59 with images), while images offer only marginal improvements. In contrast, the E2E OmniFusion model benefits substantially from both gated fusion (outperforming mid-only fusion) and image context, highlighting the importance of gating and additional contextual information. Although direct E2E translation slightly trails the best cascaded system, enabling self-cascade inference nearly closes the gap (cascaded LoRA FT: 86.59 vs. E2E gated fusion + self-cascade: 86.57). More importantly, error analysis shows that our best E2E configuration produces fewer critical and major errors (e.g., 719 vs. 751.3 major errors for the best cascaded system). These results demonstrate that OmniFusion with image context and self-cascade inference matches fine-tuned cascaded systems in offline ST quality while producing fewer severe translation errors.
4.3 Text-Image Translation
@@
Caption translation relies heavily on visual context, where cascaded systems often fail due to limited disambiguation. To evaluate OmniFusion’s ability to leverage images, we report COMET scores on the CoMMuTE test set across multiple language pairs in Table˜2. Omni 2.5B exhibits the weakest multilingual performance, motivating its fusion with a stronger translation backbone. SeedX 7B achieves reasonable scores but consistently underperforms vision-aware models, highlighting its limitations without image context. Among multimodal systems, both Mid Fusion and Gated Fusion variants of OmniFusion 14B reach near state-of-the-art performance, matching or surpassing vision-only models such as ZeroMMT 3.3B (Futeral et al., 2025) and TowerVision 9B (Viveiros et al., 2025) on most language directions.
Unlike in ST, Mid Fusion slightly outperforms Gated Fusion on CoMMuTE, suggesting that when images provide auxiliary rather than primary information, mid-layer visual integration is sufficient. Notably, unlike TowerVision, which relies on large-scale multimodal pretraining from scratch, OmniFusion builds on strong pretrained components, offering a more data- and training-efficient approach without sacrificing quality. These results highlight the context-awareness of our E2E design for visually grounded translation.
4.4 Layer Contribution - Training
Model Mode Fusion Layers en $\rightarrow$ de cn $\rightarrow$ de jp $\rightarrow$ ru fr $\rightarrow$ ko ur $\rightarrow$ cn
Chrf Comet-22 Chrf Comet-22 Chrf Comet-22 Chrf Comet-22 Chrf Comet-22
Tower-Spire Direct 56.9 82.54 N/A N/A N/A N/A N/A N/A N/A N/A
Tower-Spire Self-Cascade 61.9 85.02 N/A N/A N/A N/A N/A N/A N/A N/A
SeamlessM4T v2 Direct 62.4 85.60 35.8 70.22 36.9 71.77 20.8 74.12 13.5 61.11
SeamlessM4T v2 Self-Cascade 63.3 85.81 36.9 71.34 38.2 69.85 21.0 75.22 13.5 63.75
Our Models - OmniFusion
Attention Direct All 60.9 84.02 37.4 72.96 26.6 65.62 21.2 75.36 4.3 43.66
Attention Direct 3 61.8 86.95 39.4 74.86 27.6 68.81 21.9 78.94 5.5 46.78
Gated Direct All 62.3 85.37 39.6 75.91 28.8 69.33 21.8 77.50 5.0 45.78
Gated Direct 3 64.0 87.12 44.5 81.82 37.2 81.95 25.5 82.74 12.5 65.60
Table 3: Ablation study of fusion architecture with non-english zero-shot source languages. Tower-Spire only supports English audio.
We use only the first, middle, and last layers in the gated fusion. To validate this choice, we perform ablation experiments where all layers of the MMFM are used within the gating module. In addition, we explore an attention-based fusion approach, where a learned query token attends over all layer hidden states to produce new contextualized representations. Moreover, we compare our method with other state-of-the-art models and report the results in Table 3 on the FLEURS test set to account for multilinguality.
First, we observe that the gated variants achieve the best performance across all language pairs, including challenging ones such as Urdu to Chinese. Next, using all layers in the gating module leads to performance degradation, which becomes even more pronounced with attention-based fusion. We attribute this primarily to the increased noise introduced when incorporating all layers. Further, attention-based fusion strategies fail compared with gating due to training/test mismatch, especially for non english source languages. In comparison, our proposed approach introduces minimal number of parameters, avoiding reliance on large datasets while still achieving competitive performance.
4.5 Layer Contribution - Inference
@@
We introduced a gating module (Section˜2.1) to fuse information from multiple MMFM layers. To analyze its behavior, we perform an ablation study by examining the average gate values assigned to the first, middle, and last MMFM layers, noting that gate weights sum to 1 at each timestep. We run inference on 100 MCIF examples under speech-only and speech+image conditions, and on CoMMuTE for image-only input. The resulting average gate values are shown in Figure˜4. Across settings, first and middle layers receive substantially higher weights, while the last layer contributes marginally. This indicates that most transferable multimodal information is captured in earlier MMFM layers.
Inference-time Gate Force Speech Speech + Image
First 55.22 63.02
Middle Halluc. + ASR Halluc. + ASR
Last EOS EOS
89.90 90.05
Table 4: Inference time control of the gating with forcing first, middle or last layer. We report XCOMET-XL scores on the MCIF en $\rightarrow$ de test set.
Next, we conduct another ablation by performing inference-time control in the gating module. We ignore the gate’s predictions and instead force the model to use only the first, middle, or last layer, and report the results in Table 4. Forcing the first layer degrades translation quality but does not collapse the system and even improves performance when images are present. In contrast, forcing the middle or last layer leads to failure modes. With the middle layer, the system often generates the transcript along with additional noise, while forcing the last layer causes the model to predict only the eos token. We hypothesize that the first layer provides reasonably aligned multimodal embeddings, whereas the middle layer contains more processed representations that become ineffective due to a training–test mismatch. Notably, training with only a single layer still produces workable systems, although using the last layer yields the lowest performance and weakest zero-shot transfer, as shown in Table 5.
@@ 5 Related Work
Multimodal MT: Translating captions has been widely explored over the past decade (Specia et al., 2016; Elliott et al., 2017; Barrault et al., 2018). Several works extend NMT models to process images, either by training from scratch (Vijayan et al., 2024) or by augmenting pretrained models such as NLLB (Costa-Jussà et al., 2022; Futeral et al., 2025). Recently, and in parallel with our work, Viveiros et al. (2025) extend Tower with image and video modalities through a computationally expensive multi-stage training pipeline. Beyond vision, another line of research explores incorporating audio into translation LLMs (Ambilduke et al., 2025). A smaller body of work investigates the joint use of vision and audio, such as Sinhamahapatra and Niehues (2025), which leverages slides and speech for ASR. Recent studies also explore translating with video (Anwar et al., 2023; Choi et al., 2024). In contrast, our approach focuses on jointly leveraging images and audio while remaining efficient for data and latency.
Simultaneous ST: Real-world scenarios require real-time ST without waiting for the full utterance. However, most research trains models in offline settings and later adapts them to SimulST. Such adaptation typically relies on predefined or learned translation policies, such as wait-k (Ma et al., 2019; Elbayad et al., 2020), which assumes a fixed token delay before committing to output. Other approaches include Local Agreement (Liu et al., 2020), which selects the most stable prefix across consecutive read steps, and attention-based heuristics (Papi et al., 2023). More recently, several works explore leveraging LLMs for SimulST policy learning (Koshkin et al., 2024b, a; Guo et al., 2025; Fu et al., 2025). In this work, we adopt Local Agreement to evaluate our approach and compare against cascaded baselines. Our method is, however, compatible with other SimulST policies given its E2E nature.
Modality Fusion: A key challenge is how to effectively integrate speech and image modalities into text-centric LLMs. Some approaches (Abdin et al., 2024; Xu et al., 2025; Team et al., 2025) process each modality with separate encoders and merge them within the LLM. In contrast, models such as (Ye et al., 2025) enable early interaction between vision and audio modalities, facilitating better cross-modal alignment. Beyond the timing of such interaction, recent works also investigates the design of modality projectors (Verdini et al., 2024; Gaido et al., 2024b) and the use of discrete tokens (Zhan et al., 2024; Li et al., 2025), balancing between latency and translation quality.
@@ 6 Conclusion
We propose a novel approach for fusing an MMFM with a translation LLM to enable high-quality, multilingual, multimodal, and simultaneous translation. We validate its effectiveness across three scenarios. First, the fused model achieves lower latency while maintaining strong translation quality in SimulST, even with partial source audio. Second, it produces fewer major and critical errors. Third, it effectively exploits visual context for both speech and text translation, yielding consistent gains in multimodal settings. Our layer-level analysis further shows that the first and middle MMFM layers contribute most to semantic multimodal representations.
@@
Our training data is primarily English-centric, which may limit performance for other source languages. Additionally, while we evaluate multimodal translation on established benchmarks, broader human evaluation would help better assess the model’s ability to utilize visual context. Future work will explore multilingual multimodal training and larger-scale human evaluation Another limitation of our work is that OmniFusion is trained with fixed task-specific prompts across the three supported translation settings, and therefore does not exhibit general instruction-following capabilities. It remains unclear whether our fusion strategy can be extended to support more flexible, instruction-driven multimodal behavior.